Can anyone please help me solve this problem? Any help would be appreciated! 50 students were given a survey on jellybeans. 30 like red, 29 like green and 17 like both red and green. How many of the students like neither red or green?


Answer by stanbon(75887)

50 students were given a survey on jellybeans. 30 like red, 29 like green and 17 like both red and green.
------------------------------------
Draw a Venn diagram.
Draw two overlappiing circles inside a rectangle.
Lable one circle "red", the other "green".
put 17 in the intersection of the circles.
---
put 30-17 = 13 in the red circle but not in the intersection.
---
put 29-17 = 12 in the green circle but not in the intersection.
---
How many of the students like neither red or green?
Total # in the rectangle is 50.
Total # in the two circles is 13 + 17 + 12 = 42
------
So, total # in the rectangle and not in the circles is 8.
==========================================================
Cheers,
Stan H.
=======================